30ml of a H2O2 solution after acidification required 30 ml of 0.1 N KMnO4 solution for complete oxidation calculate the percentage of H2O2. 
Pls explain the procedure (ans is 0.17%)

Grade:11

2 Answers

User
6 years ago
Equivalent weight of H2O2= equivalent weight of KMnO4Equivalent weight of H2O2=0.03×0.1=0.003Wt. of H2O2= 0.003×34/2(n factor of H2O2=2) = 0.051gMass by vol. %= 0.051g/30ml ×100= 0.17%
User
6 years ago
Gram equivalent of H2O2=Gram equivalent of KMnO4Gram equivalent of H2O2=0.03×0.1=0.003Wt. of H2O2=0.003×34/2(n factor of H2O2=2)=0.051gMass by vol.%=0.051g/30ml=0.17% By Shivam Aditya